PROBLEM 1 [points: 20; cach sub problem is 5 pts] Note that there is no universal set present in formulas, U there is a sign expressing union operation on sets. Given set expressions, indicate which areas should be shaded, if any, on given Venn diagrams. Be careful about the order of operations.
media%2Feb8%2Feb88ac7e-eafd-41a7-9a98-97
Answer #1

ANSWER :

(c) A∩B∩(C∪B)'

First we will solve (C∪B)'
C∪B = {1,3,4,7}∪{1,2,4,6} = {1,2,3,4,6,7}
(C∪B)' = U - (C∪B ) = {5,8}
B∩(C∪B)' = {Φ}
A∩B∩(C∪B)' = {Φ}

(d) A∩B∩C'∪B
C' = U - C = {2,5,6,8}
C'∪B = {1,2,4,5,6,8}
B∩C'∪B = {1,2,4,6}
A∩B∩C'∪B = {1,2}